find three consecutive odd integers such that the squar of the thid integer plus the product of the other two integers is 268. Answer by nerdybill(7384) 
 find three consecutive odd integers such that the squar of the thid integer plus the product of the other two integers is 268.
.
Let x = first consecutive odd integer
then
x+2 = second consecutive odd integer
x+4 = third consecutive odd integer
.
From:"the square of the third integer plus the product of the other two integers is 268" we get:
(x+4)^2 + x(x+2) = 268
x^2+8x+16 + x^2+2x = 268
2x^2 + 10x + 16 = 268
x^2 + 5x + 8 = 134
x^2 + 5x - 126 = 0
(x-9)(x+14) = 0
.
x = {9, -14}
.
We can toss out the -14, leaving us with:
x=9
.
Solution:
9, 11, 13